A piece of unknown metal of mass 0.157 kg loses 1090 joules of heat when cooled from 175 degrees celsius to 25 degrees celsius. What is the specific heat capacity value of this sample?


1
Expert's answer
2021-12-14T04:52:34-0500

The equation that relates heat "q" to specific heat "c," mass "m," and temperature change "\Delta T" is shown below


"q=cm\Delta T"

Solve for "c"


"c=\dfrac{q}{m\Delta T}"

"c=\dfrac{-1090J}{0.157kg(25\degree C-175\degree C)}\approx46.2845\dfrac{J}{kg\cdot\degree C}"

"=0.0462845\dfrac{J}{g\cdot\degree C}"